OpenStudy (anonymous):

2y-4(-0.3333333333)=10

OpenStudy (anonymous):

2y-4(-1/3)=10

OpenStudy (anonymous):

2y+4/3=10 can you alve from here?

OpenStudy (anonymous):

solve*

OpenStudy (anonymous):

how did you come up with that from-0.3333333333?

OpenStudy (anonymous):

just wondering

OpenStudy (anonymous):

2y = 10-4/3 = 26/30 = 13/15.... 2y=13/15, y=(13/15)/2 = 13/30

OpenStudy (anonymous):

I came up with this answer y = 4.933333333

OpenStudy (anonymous):

if ever you have a decimal, like 0.3333 to convert it into a fraction, all you have to do multiply and divide to make it of the form 3333/10000, (the number of zeroes and the number of digts after the decimal point shall always match in such cases), and 33/100 can approximately be simplified to 1/3

OpenStudy (anonymous):

i hate algebra

OpenStudy (anonymous):

so is that like1÷3?